### v2

@@ 3.2.4 Data Collection Challenges and Solutions
Lin et al. lin2018towards presented a Human-centred Agile Software Engineering (HASE) platform. The HASE platform is an online Agile Project Management (APM) tool allowing unobtrusive data collection during the normal ASD process. It collects data on software engineering activities from the participants, such as task allocation, collaboration, and mood stability, without requiring additional effort from the participants. This approach ensures that the data collection process does not disrupt the regular workflow of the software engineering teams while still providing accurate and objective data for assessing software engineering skills. Svensson et al. svensson2019unfulfilled underscored that future research should develop novel automated methods for collecting, analyzing, and visualizing data to enhance existing agile decision-making processes by associating pertinent data with specific decision scenarios. Upender upender2005staying proposed a three-year project to build a centralized data management system that addresses the needs of electronically collecting and integrating data with a gradual adoption and adaptation of Scrum and XP (Extreme Programming) practices. They emphasized a ”just in time design” approach, which meant they approached the design phase with a flexible mindset. They used an evolutionary design process to handle ambiguous and conflicting requirements. Initially, they built a basic version of the clinical application to ensure it had essential functions, leaving the implementation of more advanced features for future updates. At the time the study was written, they were two years into the project.
Data Sharing and Collaboration: Fabijan et al. fabijan2016lack noted the lack of customer and product data sharing among agile team members. This leads to difficulties in accessing data collected by others in different development phases and fragmented collection and storage of data. Kaur et al. kaur2020dialogue pinpointed concurrent access as a significant challenge, leading to unnecessary rework due to human error and inefficient communication methods. Barbala et al. barbala2023data discussed the challenge of the uncertainty surrounding what data agile product teams are allowed to gather.
@@
Informative Data Collection Challenges: Dos Santos Júnior et al. dos2021ontology delved into difficulty in accessing, integrating, analyzing, and viewing data handled by heterogeneous applications, which often adopt different semantics (the same information item is given divergent interpretations), posing a barrier to integrated data usage. Matthies matthies2020playing discussed the challenge of relying on subjective data gathered from team members’ perceptions and experiences during Retrospective meetings. While readily available and relevant to team satisfaction, this subjective data may not provide a comprehensive or objective view of the team’s performance and areas for improvement. Martínez-Fernández et al. martinez2019continuously addressed the data collection challenge of obtaining informative user data. Dharmapal and Sikamani dharmapal2016big underscored the significance of collecting data from suitable sources and filtering out unwanted data in the development phase.
To address the informative data collection challenges, Dos Santos Júnior et al. dos2021ontology proposed a solution “Development of ontology-based approach Immigrant” discussed in Section 3.2.3. This solution not only resolved the semantic heterogeneity data integration challenge but also assisted in collecting informative data that can be used for data-driven decision-making in agile software development. Matthies matthies2020playing recommended integrating development project data and insights from software repository mining into Retrospective agendas. By employing project data from the last iteration in Retrospective meetings, teams can gain additional insights based on their team-specific data. This approach aims to provide a more thorough overview of the team’s state, including facts and feelings, which can lead to better results in Retrospectives. Project data measurements can help track progress on common agile and software engineering challenges, enabling teams to identify and tackle improvement actions based on empirical project evidence rather than solely on anecdotal experiences.
@@
Martínez-Fernández et al. martinez2019continuously proposed using the Q-Rapids tool for acquiring valuable information for users. The Q-Rapids tool is designed to gather diverse data related to software system development and usage. These data are then structured within a Quality Model (QM) to analyze aggregated quality-related strategic indicators. These indicators are made available to decision-makers through a multi-dimensional and navigational dashboard employed during ASD events like sprint planning or daily stand-up meetings. The Q-Rapids solution was implemented and used in a real-world setting. Martínez-Fernández et al. martinez2019continuously described a case study across four companies involving 26 practitioners to investigate the integration of a Quality Model within the Q-Rapids software analytics tool. Dharmapal and Sikamani dharmapal2016big proposed utilizing a combination of agile methodology principles and the expertise of a Data Analyst. During the planning stage, the Product Owner and the Data Analyst collaborate to select suitable Big Data Analytics tools and strategize the data source collection. This approach allows for incremental data collection in response to specific needs. In the development phase, the Data Analyst is pivotal in analyzing data gathered from various sources. They integrate data from diverse origins into a unified dataset, merge information based on shared attributes, and present it in a consolidated format. An essential aspect of their role is removing irrelevant data, employing an elimination technique that saves time and resources by excluding unnecessary data from processing.
@@
Comprehensive Data Collection: Das et al. das2015towards delved into the challenges associated with collecting and storing substantial amounts of data for specific research purposes, mainly when the data is non-traditional, or the dataset is immense. Huang et al. huang2021leveraging focused on efficiently ingesting and governing large volumes of heterogeneous data from various sources in various formats. Fagarasan et al. fagarasan2023integrating stressed the importance of systematic and comprehensive data collection, beginning at the lowest level of project management. Olsson olsson2018challenges discussed that traditional data collection methods, such as surveys and interviews, are insufficient for continuous experimentation.
@@
To address the comprehensive data collection challenges, Das et al. das2015towards advocated for a two-fold approach: first, the creation of a generalized dataset aimed at expanding data coverage (not volume) by incorporating additional attributes. This approach ensures that a single data extraction step can serve multiple projects, with adjustments made in subsequent workflows to cater to specific research task requirements. Second, they emphasized standardized data processing, which involves abstracting standard data procedures, such as noise reduction and language standardization, into reusable solutions and tools applicable across various data-driven research endeavors. It makes Datto systematically collect and analyze. Huang et al. huang2021leveraging The solution they proposed was to design a big data stack with scalable storage technologies and a flexible architecture. This stack included a data lake for raw data storage and integrated relational databases for real-time applications. Apache Spark was used as a distributed data computation engine to enable parallel processing for analytics tasks. The stack’s design allowed for swift data analysis and agile development of data products, reducing the time required to answer analytical questions and develop applications. Fagarasan et al. fagarasan2023integrating proposed using a market-available Project Management Information System (PMIS) to collect and analyse data systematically. This system enables the automated collection and reporting of key performance indicators, crucial for monitoring project performance and ensuring cost-effective execution. The PMIS is a foundation for organizing projects into a measurable portfolio, with performance metrics deeply embedded in the software development workflow. Olsson olsson2018challenges proposed implementing mechanisms for automated data collection directly from the software in use, particularly in the post-deployment stage. This approach allows for the collection of real-time and actionable data that can be used for analysis to improve current products and inform the development of future ones, aligning with legal requirements for data utilization and user consent.
@@ 3.2.5 Data Quality Challenges and Solutions
Table 4: Summary of Data Quality Challenges and Solutions
Data Quality Challenge Data Quality Solution Solution Status Ref.
Ensuring Data Accuracy and Consistency Across Varied Sources Solution referenced in Section 3.2.3 Implemented rosenkranz2017supporting
Solution referenced in Section 3.2.3 Implemented harper2014agile
Solution referenced in Section 3.2.3 Implemented little2004adaptive
Solution referenced in Section 3.2.4 Implemented pater2018advancing
Solution referenced in Section 3.2.3 Implemented kannan2017rapid
Missing Quality Data Utilizing a quality-aware rapid software Implemented franch2019quality
Solution referenced in Section 3.2.4 Proposed lehtonen2017visualizations
Solution referenced in Section 3.2.4 Proposed martinez2019continuously
Inadequate Data Quality Management Utilizing earlier test-driven approaches Proposed ambler2008gets
Solution referenced in Section 3.2.3 Implemented hofer2020new
Solution referenced in Section 3.2.3 Implemented svensson2019unfulfilled
Data Quality Standardization Solution referenced in Section 3.2.4 Implemented spengler2020enabling
Solution referenced in Section 3.2.4 Proposed batarseh2018predicting
@@
Ensuring Data Accuracy and Consistency Across Varied Sources: Rosenkranz et al. rosenkranz2017supporting highlighted the importance of data quality (defined in their study as completeness, ambiguity, meaningfulness, and correctness) in application systems and project performance in IT or Data Warehouse (DW) projects. The challenge lies in ensuring the accuracy and consistency of data from disparate sources. Harper and Dagnino harper2014agile also emphasized the need to merge and clean datasets from different sources. The challenge here is ensuring the accuracy and consistency of data for analysis. Little et al. little2004adaptive mentioned that the products are released regularly, with release cycles ranging from 3 to 18 months. This regularity poses a challenge in managing data and ensuring consistency across different product versions. Pater et al. pater2018advancing underscored the necessity to eradicate data redundancy and business process fragmentation in the context of IT system modernization. The challenge lies in aligning the diverse IT needs of users with the broader organizational data requirements. Kannan et al. kannan2017rapid highlighted the challenge of ensuring the completeness, accuracy, and internal consistency of the critical few EHR fields essential for registry inclusion and eCQM (electronic Clinical Quality Measure) calculation.
@@
To address the data accuracy and consistency challenges across varied sources: The solutions are the same as those of the data integration and collection (as those challenges co-occur together). We discussed those solutions in detail in Sections 3.2.3 and 3.2.4.
@@
Missing Quality Data: Franch et al. franch2019quality emphasized the issue when necessary quality data is unavailable (i.e., inaccessible, incomplete, inconsistent, and incorrect data) due to the absence of suitable data collection tools or processes or inaccessible data storage systems. For instance, SonarQube (mentioned as an existing tool that focuses on product quality or continuous integration) does not furnish raw data for specific base metrics. Lehtonen et al. lehtonen2017visualizations highlighted potential data quality problems in software engineering data, including noise, outliers, low precision, missing values, coverage errors, and clones. An example given is the potential for inaccurate data due to developers forgetting to update the issue management system task state when development work starts or ends, leading to inaccurate timestamps. Mart’ınez-Fern’andez martinez2019continuously discussed the need for transparency and clarity on the raw data used to compute factors and indicators in the software analytics tool (Q-Rapids tool). End-users like product owners may require a clear understanding of the raw data used to derive specific values and the associated decision-making processes.
@@
To address the missing data Franch et al. franch2019quality discussed the Q-Rapids project, mentioned as an initiative aimed at developing a quality-aware rapid software development methodology. The Q-Rapids project addressed data quality challenges by selecting and configuring tailored data collection tools aligned with business objectives, specifying precision requirements, implementing real-time checks to prevent errors, and integrating diverse data sources while transforming them for compatibility with specific analysis methods and tools. The solution was implemented, evaluated, and refined through three releases deployed by the four industry partners in the project. Note that the solutions suggested to resolve the remaining data quality challenges, as discussed by Lehtonen et al. lehtonen2017visualizations and Mart’ınez-Fern’andez martinez2019continuously , are the same as those of data collection (as they co-occur together). We discussed those solutions in detail in Section 3.2.4.
@@
Inadequate Data Quality Management: Ambler ambler2008gets highlighted that data quality issues cost US organizations an estimated 600 billion dollars annually (according to Data Warehouse Institute), stemming from the use of traditional data management approaches to ensure data quality within organizations, according to the Data Warehouse Institute as indicated by a survey conducted by Ambler ambler2008gets . Svensson et al. svensson2019unfulfilled emphasized that the quality of the data and the processing techniques and tools used to handle it will impact the quality of decisions made using Data-Driven Decision Making (DDDM). Hofer et al. hofer2020new discussed the data quality challenge of ensuring the ”fitness for use” of data, which is often neglected or delayed in the software engineering process until the end-user evaluation phase, known as the “point-of-truth”. This delay impacts the cost-effectiveness of data quality management and contradicts agile principles that emphasize early and continuous delivery of valuable software.
@@
Data Quality Standardization: Spengler et al. spengler2020enabling mentioned the need for significant data restructuring and cleansing due to heterogeneous data to ensure suitability for proper integration into clinical data warehousing solutions. Batarseh and Gonzalez batarseh2018predicting highlighted the need to enhance dataset quality, standardize variables, and eliminate unnecessary data. The challenge here involves dealing with poor-quality data and standardizing data for analysis.

### v4

@@ 3.2.5 Data Collection Challenges and Solutions
Lin et al. lin2018towards presented a Human-centred Agile Software Engineering (HASE) platform. The HASE platform is an online Agile Project Management (APM) tool allowing unobtrusive data collection during the normal ASD process. It collects data on software engineering activities from the participants, such as task allocation, collaboration, and mood stability, without requiring additional effort from the participants. This approach ensures that the data collection process does not disrupt the regular workflow of the software engineering teams while still providing accurate and objective data for assessing software engineering skills. Svensson et al. svensson2019unfulfilled underscored that future research should develop novel automated methods for collecting, analyzing, and visualizing data to enhance existing agile decision-making processes by associating pertinent data with specific scenarios. Upender upender2005staying proposed a three-year project to build a centralized data management system that addresses the needs of electronically collecting and integrating data with a gradual adoption and adaptation of Scrum and XP (Extreme Programming) practices. They emphasized a ”just in time design” approach, which meant they approached the design phase with a flexible mindset. They used an evolutionary design process to handle ambiguous and conflicting requirements. Initially, they built a basic version of the clinical application to ensure it had essential functions, leaving the implementation of more advanced features for future updates. At the time the study was written, they were two years into the project.
Data Sharing and Collaboration: This challenge concerns ensuring data is accessible and shared seamlessly across agile teams. Limited access to data and fragmented storage often hinder collaboration and transparency. Fabijan et al. fabijan2016lack noted the lack of customer and product data sharing among agile team members. This leads to difficulties in accessing data collected by others in different development phases and fragmented collection and storage of data. Kaur et al. kaur2020dialogue pinpointed concurrent access as a significant challenge, leading to unnecessary rework due to human error and inefficient communication methods. Barbala et al. barbala2023data discussed the challenge of the uncertainty surrounding what data agile product teams are allowed to gather.
@@
Informative Data Collection: This challenge refers to the difficulties agile teams face in gathering meaningful and actionable data for decision-making. While data is abundant, its quality and relevance often fall short of agile requirements. Dos Santos Júnior et al. dos2021ontology delved into difficulty in accessing, integrating, analyzing, and viewing data handled by heterogeneous applications, which often adopt different semantics (the same information item is given divergent interpretations), posing a barrier to integrated data usage. Matthies matthies2020playing discussed the challenge of relying on subjective data gathered from team members’ perceptions and experiences during Retrospective meetings. While readily available and relevant to team satisfaction, this subjective data may not provide a comprehensive or objective view of the team’s performance and areas for improvement. Martínez-Fernández et al. martinez2019continuously addressed the data collection challenge of obtaining informative user data. Dharmapal and Sikamani dharmapal2016big underscored the significance of collecting data from suitable sources and filtering out unwanted data in the development phase.
@@
To address the informative data collection challenges, Dos Santos Júnior et al. dos2021ontology proposed a solution “Development of ontology-based approach Immigrant” discussed in Section 3.2.4. This solution resolved the semantic heterogeneity data integration challenge and assisted in collecting informative data that can be used for data-driven decision-making in agile software development. Matthies matthies2020playing recommended integrating development project data and insights from software repository mining into Retrospective agendas. By employing project data from the last iteration in Retrospective meetings, teams can gain additional insights based on their team-specific data. This approach aims to provide a more thorough overview of the team’s state, including facts and feelings, which can lead to better results in Retrospectives. Project data measurements can help track progress on common agile and software engineering challenges, enabling teams to identify and tackle improvement actions based on empirical project evidence rather than solely on anecdotal experiences. Martínez-Fernández et al. martinez2019continuously proposed using the Q-Rapids tool for acquiring valuable information for users. The Q-Rapids tool is designed to gather diverse data related to software system development and usage. These data are then structured within a Quality Model (QM) to analyze aggregated quality-related strategic indicators. These indicators are made available to decision-makers through a multi-dimensional and navigational dashboard employed during ASD events like sprint planning or daily stand-up meetings. The Q-Rapids solution was implemented and used in a real-world setting. Martínez-Fernández et al. martinez2019continuously described a case study across four companies involving 26 practitioners to investigate the integration of a Quality Model within the Q-Rapids software analytics tool. Dharmapal and Sikamani dharmapal2016big proposed utilizing a combination of agile methodology principles and the expertise of a Data Analyst. During the planning stage, the Product Owner and the Data Analyst collaborate to select suitable Big Data Analytics tools and strategize the data source collection. This approach allows for incremental data collection in response to specific needs. In the development phase, the Data Analyst is pivotal in analyzing data gathered from various sources. They integrate data from diverse origins into a unified dataset, merge information based on shared attributes, and present it in a consolidated format. An essential aspect of their role is removing irrelevant data, employing an elimination technique that saves time and resources by excluding unnecessary data from processing.
@@
Comprehensive Data Collection: This involves collecting a wide range of data from various sources to support agile decision-making. The challenge lies in processing and integrating diverse datasets, especially in environments with large-scale or unstructured data. Das et al. das2015towards delved into the challenges associated with collecting and storing substantial amounts of data for specific research purposes, mainly when the data is non-traditional or the dataset is immense. Huang et al. huang2021leveraging focused on efficiently ingesting and governing large volumes of heterogeneous data from various sources in various formats. Fagarasan et al. fagarasan2023integrating stressed the importance of systematic and comprehensive data collection, beginning at the lowest level of project management. Olsson olsson2018challenges discussed that traditional data collection methods, such as surveys and interviews, are insufficient for continuous experimentation.
@@
To address the comprehensive data collection challenges, Das et al. das2015towards advocated for a two-fold approach: first, the creation of a generalized dataset aimed at expanding data coverage (not volume) by incorporating additional attributes. This approach ensures that a single data extraction step can serve multiple projects, with adjustments made in subsequent workflows to cater to specific research task requirements. Second, they emphasized standardized data processing, which involves abstracting standard data procedures, such as noise reduction and language standardization, into reusable solutions and tools applicable across various data-driven research endeavors. It makes Datto systematically collect and analyze. Huang et al. huang2021leveraging The solution they proposed was to design a big data stack with scalable storage technologies and a flexible architecture. This stack included a data lake for raw data storage and integrated relational databases for real-time applications. Apache Spark was used as a distributed data computation engine to enable parallel processing for analytics tasks. The stack’s design allowed for swift data analysis and agile development of data products, reducing the time required to answer analytical questions and develop applications. Fagarasan et al. fagarasan2023integrating proposed using a market-available Project Management Information System (PMIS) to collect and analyze data systematically. This system enables the automated collection and reporting of key performance indicators, which are crucial for monitoring project performance and ensuring cost-effective execution. The PMIS is a foundation for organizing projects into a measurable portfolio, with performance metrics deeply embedded in the software development workflow. Olsson olsson2018challenges proposed implementing mechanisms for automated data collection directly from the software in use, particularly in the post-deployment stage. This approach allows for the collection of real-time and actionable data that can be used for analysis to improve current products and inform the development of future ones, aligning with legal requirements for data utilization and user consent.
@@ 3.2.6 Data Quality Challenges and Solutions
Table 5: Summary of Data Quality Challenges and Solutions Identified by the SLR.
Data Quality Challenge Solution Implementation Status Ref. Ensuring Data Accuracy and Consistency Across Varied Sources Ref in Sec. 3.2.4 Implemented rosenkranz2017supporting Ref in Sec. 3.2.4 Implemented harper2014agile Ref in Sec. 3.2.4 Implemented little2004adaptive Ref in Sec. 3.2.5 Implemented pater2018advancing Ref in Sec. 3.2.4 Implemented kannan2017rapid Missing Quality Data Utilizing a quality -aware models Implemented franch2019quality Ref in Sec. 3.2.5 Proposed lehtonen2017visualizations Ref in Sec. 3.2.5 Proposed martinez2019continuously Inadequate Data Quality Management Using test-driven approaches Proposed ambler2008gets Ref in Sec. 3.2.4 Implemented hofer2020new Ref in Sec. 3.2.4 Implemented svensson2019unfulfilled Data Quality Standardization Ref in Sec. 3.2.5 Implemented spengler2020enabling Ref in Sec. 3.2.5 Proposed batarseh2018predicting
@@
Ensuring Data Accuracy and Consistency Across Varied Sources: This challenge arises when data from multiple systems and sources must be aligned to ensure reliability and usability. In agile environments, where frequent updates and dynamic data integration are common, maintaining accuracy and consistency is critical to avoid errors and inefficiencies. Rosenkranz et al. rosenkranz2017supporting highlighted the importance of data quality (defined in their study as completeness, ambiguity, meaningfulness, and correctness) in application systems and project performance in IT or Data Warehouse (DW) projects. The challenge lies in ensuring the accuracy and consistency of data from disparate sources. Harper and Dagnino harper2014agile also emphasized the need to merge and clean datasets from different sources. The challenge here is ensuring the accuracy and consistency of data for analysis. Little et al. little2004adaptive mentioned that the products are released regularly, with release cycles ranging from 3 to 18 months. This regularity poses a challenge in managing data and ensuring consistency across different product versions. Pater et al. pater2018advancing underscored the necessity to eradicate data redundancy and business process fragmentation in the context of IT system modernization. The challenge lies in aligning the diverse IT needs of users with the broader organizational data requirements. Kannan et al. kannan2017rapid highlighted the challenge of ensuring the completeness, accuracy, and internal consistency of the critical few EHR fields essential for registry inclusion and eCQM (electronic Clinical Quality Measure) calculation.
@@
To address the data accuracy and consistency challenges across varied sources: The solutions are the same as those of the data integration and collection (as those challenges co-occur together). We discussed those solutions in detail in Sections 3.2.4 and 3.2.5.
@@
Missing Quality Data: This challenge pertains to the absence of complete, accurate, or usable data, which can hinder agile teams’ ability to make informed decisions. The lack of quality data in agile development often stems from incomplete data collection processes or insufficient tools. Franch et al. franch2019quality emphasized the issue when necessary quality data is unavailable (i.e., inaccessible, incomplete, inconsistent, and incorrect data) due to the absence of suitable data collection tools or processes or inaccessible data storage systems. For instance, SonarQube (mentioned as an existing tool that focuses on product quality or continuous integration) does not furnish raw data for specific base metrics. Lehtonen et al. lehtonen2017visualizations highlighted potential data quality problems in software engineering data, including noise, outliers, low precision, missing values, coverage errors, and clones. An example given is the potential for inaccurate data due to developers forgetting to update the issue management system task state when development work starts or ends, leading to inaccurate timestamps. Martínez-Fernández et al. martinez2019continuously discussed the need for transparency and clarity on the raw data used to compute factors and indicators in the software analytics tool (Q-Rapids tool). End-users like product owners may require a clear understanding of the raw data used to derive specific values and the associated decision-making processes.
@@
To address the missing data Franch et al. franch2019quality discussed the Q-Rapids project, mentioned as an initiative aimed at developing a quality-aware rapid software development methodology. The Q-Rapids project addressed data quality challenges by selecting and configuring tailored data collection tools aligned with business objectives, specifying precision requirements, implementing real-time checks to prevent errors, and integrating diverse data sources while transforming them for compatibility with specific analysis methods and tools. The solution was implemented, evaluated, and refined through three releases deployed by the four industry partners in the project. Note that the solutions suggested to resolve the remaining data quality challenges, as discussed by Lehtonen et al. lehtonen2017visualizations and Martínez-Fernández et al. martinez2019continuously , are the same as those of data collection (as they co-occur together). We discussed those solutions in detail in Section 3.2.5.
@@
Inadequate Data Quality Management: This refers to the failure to implement effective processes for monitoring, maintaining, and improving data quality throughout agile projects. Agile workflows often emphasize rapid delivery, which can lead to delayed focus on data quality until critical phases. Ambler ambler2008gets highlighted that data quality issues cost US organizations an estimated 600 billion dollars annually (according to Data Warehouse Institute), stemming from the use of traditional data management approaches to ensure data quality within organizations, according to the Data Warehouse Institute as indicated by a survey conducted by Ambler ambler2008gets . Svensson et al. svensson2019unfulfilled emphasized that the quality of the data and the processing techniques and tools used to handle it will impact the quality of decisions made using Data-Driven Decision Making (DDDM). Hofer et al. hofer2020new discussed the data quality challenge of ensuring the ”fitness for use” of data, which is often neglected or delayed in the software engineering process until the end-user evaluation phase, known as the “point-of-truth”. This delay impacts the cost-effectiveness of data quality management and contradicts agile principles that emphasize early and continuous delivery of valuable software.
@@
Data Quality Standardization: This involves ensuring that data adheres to consistent formats and standards to support integration and analysis. Achieving standardization is critical in agile environments, where data is sourced from diverse systems and rapidly evolving requirements. Spengler et al. spengler2020enabling mentioned the need for significant data restructuring and cleansing due to heterogeneous data to ensure suitability for proper integration into clinical data warehousing solutions. Batarseh and Gonzalez batarseh2018predicting highlighted the need to enhance dataset quality, standardize variables, and eliminate unnecessary data. The challenge here involves dealing with poor-quality data and standardizing data for analysis.